# Deeper Insights into Deep Graph Convolutional Networks: Stability and Generalization

### v1

@@ Appendix A Proofs
    $\|\mathbf{A}_{1}\mathbf{A}_{2}-\mathbf{A}^{\prime}_{1}\mathbf{A}^{\prime}_{2}\|_{F}\leq\|\mathbf{A}_{1}-\mathbf{A}_{1}^{\prime}\|_{F}\|\mathbf{A}_{2}\|_{2}+\|\mathbf{A}_{2}-\mathbf{A}_{2}^{\prime}\|_{F}\|\mathbf{A}^{\prime}_{1}\|_{2}.$ To show this, note that
@@
    Then, the proof is complete using the first inequality $\|\mathbf{A}_{1}\mathbf{A}_{2}\|_{F}\leq\|\mathbf{A}_{1}\|_{2}\|\mathbf{A}_{2}\|_{F}$,
@@ A.1 Gradient computation for SGD
    $\nabla_{\mathbf{w}}f(\mathbf{x}|\theta)=\nabla\sigma\big{(}\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}\big{)}\big{[}\boldsymbol{\delta}^{\top}_{\mathbf{x}}g(\mathbf{L})\mathbf{X}^{(K)}\big{]}^{\top},$ (19)
@@
    $\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta)=\big{[}g(\mathbf{L})\mathbf{X}^{(k-1)}\big{]}^{\top}\Big{(}\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\odot\mathbf{R}^{(k)}\Big{)},$ (20)
@@
    $\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k-1)}}=g(\mathbf{L})^{\top}\Big{(}\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\odot\mathbf{R}^{(k)}\Big{)}\big{[}\mathbf{W}^{(k)}\big{]}^{\top},$ (21)
@@
    $\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(K)}}=\nabla\sigma\big{(}\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}\big{)}\big{[}\boldsymbol{\delta}^{\top}_{\mathbf{x}}g(\mathbf{L})\big{]}^{\top}\mathbf{w}^{\top},$ (22)
The notation $\odot$ represents the Hadamard product of two matrices. (19) and (22) are easy to verify, while (20) and (21) are not. In the following, a detailed procedure is provided to derive (20) and (21).
@@
$\frac{\partial\mathbf{X}^{(k)}_{ij}}{\partial\mathbf{X}^{(k-1)}}$ $=\frac{\partial\sigma\big{(}\boldsymbol{\delta}_{i}^{\top}g(\mathbf{L})\mathbf{X}^{(k-1)}\mathbf{W}^{(k)}\boldsymbol{\delta}_{j}\big{)}}{\partial\mathbf{X}^{(k-1)}}$
$=\nabla\sigma\big{(}\boldsymbol{\delta}_{i}^{\top}g(\mathbf{L})\mathbf{X}^{(k-1)}\mathbf{W}^{(k)}\boldsymbol{\delta}_{j}\big{)}g(\mathbf{L})^{\top}\boldsymbol{\delta}_{i}\boldsymbol{\delta}_{j}^{\top}\big{[}\mathbf{W}^{(k)}\big{]}^{\top}.$
@@
This completes the derivation of (20) and (21).
@@
    $\|\mathbf{X}^{(k)}\|_{F}\leq B^{k}\alpha^{k}_{\sigma}C^{k}_{g}C_{\mathbf{X}}.$ (23)
@@
    $\|\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\|_{F}\leq B^{K+1-k}\alpha^{K+1-k}_{\sigma}C^{K+1-k}_{g}.$ (24)
@@
    $\big{\|}\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta)\big{\|}_{F}\leq B^{K}\alpha_{\sigma}^{K+1}C^{K+1}_{g}C_{\mathbf{X}},$ (25)
@@
    Firstly, for $k=2,3,\dots,K$,
    $\|\mathbf{X}^{(k)}\|_{F}$ $=\|\sigma\big{(}g(\mathbf{L})\mathbf{X}^{(k-1)}\mathbf{W}^{(k)}\big{)}\|_{F}$
    $\leq\alpha_{\sigma}\|g(\mathbf{L})\mathbf{X}^{(k-1)}\mathbf{W}^{(k)}\|_{F}$
    $\leq B\alpha_{\sigma}C_{g}\|\mathbf{X}^{(k-1)}\|_{F}.$
    Since $\|\mathbf{X}^{(1)}\|_{F}=\|\sigma\big{(}g(\mathbf{L})\mathbf{X}\mathbf{W}^{(1)}\big{)}\|_{F}\leq B\alpha_{\sigma}C_{g}C_{\mathbf{X}}$, we have
@@
    which completes the proof of (23).
@@
    To show (24), note that for $k=1,2,\dots,K-1$,
    $\|\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\|_{F}=$ $\|g(\mathbf{L})^{\top}\Big{(}\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k+1)}}\odot\mathbf{R}^{(k+1)}\Big{)}\big{[}\mathbf{W}^{(k+1)}\big{]}^{\top}\|_{F}$
    $\leq$ $\|g(\mathbf{L})\|_{2}\|\Big{(}\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k+1)}}\odot\mathbf{R}^{(k+1)}\Big{)}\|_{F}\|\mathbf{W}^{(k+1)}\|_{2}$
    $\leq$ $BC_{g}\|\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k+1)}}\odot\mathbf{R}^{(k+1)}\|_{F}\leq B\alpha_{\sigma}C_{g}\|\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k+1)}}\|_{F}.$
    Furthermore, since
    $\|\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(K)}}\|_{F}$ $=\|\nabla\sigma\big{(}\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}\big{)}\big{[}\boldsymbol{\delta}^{\top}_{\mathbf{x}}g(\mathbf{L})\big{]}^{\top}\mathbf{w}\|_{F}$
    $\leq B\alpha_{\sigma}C_{g},$
    then for $k=1,2,\dots,K$,
@@
    This completes the proof of (24).
@@
    To show (25), note that
    $\big{\|}\nabla_{\mathbf{w}}f(\mathbf{x}|\theta)\big{\|}_{F}=$ $\big{\|}\nabla\sigma\big{(}\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}\big{)}\big{[}g(\mathbf{L})\mathbf{X}^{(K)}\big{]}^{\top}\boldsymbol{\delta}_{\mathbf{x}}\big{\|}_{F}$
    $\leq$ $\alpha_{\sigma}\|\mathbf{X}^{(K)}\|_{F}\|\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\|_{2}\leq B^{K}\alpha_{\sigma}^{K+1}C^{K+1}_{g}C_{\mathbf{X}}.$
    Furthermore, for $k=1,2,\dots,K-1$,
    $\|\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta)\|_{F}=$ $\big{\|}\big{[}g(\mathbf{L})\mathbf{X}^{(k-1)}\big{]}^{\top}\Big{(}\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\odot\mathbf{R}^{(k)}\Big{)}\big{\|}_{F}$
    $=$ $\big{\|}g(\mathbf{L})\big{\|}_{2}\big{\|}\mathbf{X}^{(k-1)}\big{\|}_{F}\big{\|}\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\odot\mathbf{R}^{(k)}\big{\|}_{F}$
    $\leq$ $C_{g}\big{\|}\mathbf{X}^{(k-1)}\big{\|}_{F}\cdot\alpha_{\sigma}\big{\|}\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\big{\|}_{F}$
    $\leq$ $B^{K}\alpha^{K+1}_{\sigma}C^{K+1}_{g}C_{\mathbf{X}},$
    which completes the proof of (25).
A.2 Proof of Lemma 2
@@
    $\|\triangle\mathbf{X}^{(k)}\|_{F}\leq B^{k-1}\alpha^{k}_{\sigma}C^{k}_{g}C_{\mathbf{X}}\Big{(}\sum_{j=1}^{k}\|\triangle\mathbf{W}^{(j)}\|_{2}\Big{)}.$ (26)
@@
    $|f(\mathbf{x}|\theta)-f(\mathbf{x}|\theta^{\prime})|\leq B^{K}\alpha^{K+1}_{\sigma}C^{K+1}_{g}C_{\mathbf{X}}\|\triangle\theta\|_{*}.$ (27)
Proof.
To prove (26), note that for $k=1,2,\dots,K$,
$\|\triangle\mathbf{X}^{(k)}\|_{F}=$ $\|\mathbf{X}^{(k)}-\mathbf{X}^{(k)^{\prime}}\|_{F}$
$=$ $\|\sigma\big{(}g(\mathbf{L})\mathbf{X}^{(k-1)}\mathbf{W}^{(k)}\big{)}-\sigma\big{(}g(\mathbf{L})\mathbf{X}^{(k-1)^{\prime}}\mathbf{W}^{(k)^{\prime}}\big{)}\|_{F}$
$\leq$ $\alpha_{\sigma}\|g(\mathbf{L})\big{(}\mathbf{X}^{(k-1)}\mathbf{W}^{(k)}-\mathbf{X}^{(k-1)^{\prime}}\mathbf{W}^{(k)^{\prime}}\big{)}\|_{F}$
$\leq$ $\alpha_{\sigma}\|g(\mathbf{L})\|_{2}\|\mathbf{X}^{(k-1)}\mathbf{W}^{(k)}-\mathbf{X}^{(k-1)^{\prime}}\mathbf{W}^{(k)^{\prime}}\|_{F}$
$\leq$ $\alpha_{\sigma}C_{g}\Big{(}\|\mathbf{X}^{(k-1)}\|_{F}\|\triangle\mathbf{W}^{(k)}\|_{2}+\|\triangle\mathbf{X}^{(k-1)}\|_{F}\|\mathbf{W}^{(k)^{\prime}}\|_{2}\Big{)}$
$\leq$ $\alpha_{\sigma}C_{g}\Big{(}B^{k-1}\alpha^{k-1}_{\sigma}C^{k-1}_{g}C_{\mathbf{X}}\|\triangle\mathbf{W}^{(k)}\|_{2}+B\|\triangle\mathbf{X}^{(k-1)}\|_{F}\Big{)}$
$\leq$ $B^{k-1}\alpha^{k}_{\sigma}C^{k}_{g}C_{\mathbf{X}}\|\triangle\mathbf{W}^{(k)}\|_{2}+B\alpha_{\sigma}C_{g}\|\triangle\mathbf{X}^{(k-1)}\|_{F}.$
Then, since $\|\triangle\mathbf{X}^{(1)}\|_{F}\leq\alpha_{\sigma}C_{g}C_{\mathbf{X}}\|\triangle\mathbf{W}^{(1)}\|_{2}$,

### v2

@@ Acknowledgment
    Similarly, we also have $\|\mathbf{A}_{1}\mathbf{A}_{2}\|_{F}\leq\|\mathbf{A}_{1}\|_{F}\|\mathbf{A}_{2}\|_{2}$.
@@
    $\|\mathbf{A}_{1}\mathbf{A}_{2}-\mathbf{A}^{\prime}_{1}\mathbf{A}^{\prime}_{2}\|_{F}\leq\|\mathbf{A}_{1}-\mathbf{A}_{1}^{\prime}\|_{F}\|\mathbf{A}_{2}\|_{2}+\|\mathbf{A}^{\prime}_{1}\|_{F}\|\mathbf{A}_{2}-\mathbf{A}_{2}^{\prime}\|_{2}.$ To show this, note that
@@
    Then, the proof is complete using the first inequality $\|\mathbf{A}_{1}\mathbf{A}_{2}\|_{F}\leq\|\mathbf{A}_{1}\|_{F}\|\mathbf{A}_{2}\|_{2}$,
@@
    $\nabla_{\mathbf{w}}f(\mathbf{x}|\theta)=\nabla\sigma\big(\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}\big)\big[\boldsymbol{\delta}^{\top}_{\mathbf{x}}g(\mathbf{L})\mathbf{X}^{(K)}\big]^{\top},$ (A.1)
@@
    $\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta)=\big[g(\mathbf{L})\mathbf{X}^{(k-1)}\big]^{\top}\Big(\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\odot\mathbf{R}^{(k)}\Big),$ (A.2)
@@
    $\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k-1)}}=g(\mathbf{L})^{\top}\Big(\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\odot\mathbf{R}^{(k)}\Big)\big[\mathbf{W}^{(k)}\big]^{\top},$ (A.3)
@@
    $\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(K)}}=\nabla\sigma\big(\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}\big)\big[\boldsymbol{\delta}^{\top}_{\mathbf{x}}g(\mathbf{L})\big]^{\top}\mathbf{w}^{\top}.$ (A.4)
The notation $\odot$ represents the Hadamard product of two matrices. (A.1) and (A.4) are easy to verify, while (A.2) and (A.3) are not. In the following, a detailed procedure is provided to derive (A.2) and (A.3).
@@
$\frac{\partial\mathbf{X}^{(k)}_{ij}}{\partial\mathbf{X}^{(k-1)}}$ $=\frac{\partial\sigma\big(\boldsymbol{\delta}_{i}^{\top}g(\mathbf{L})\mathbf{X}^{(k-1)}\mathbf{W}^{(k)}\boldsymbol{\delta}_{j}\big)}{\partial\mathbf{X}^{(k-1)}}=\nabla\sigma\big(\boldsymbol{\delta}_{i}^{\top}g(\mathbf{L})\mathbf{X}^{(k-1)}\mathbf{W}^{(k)}\boldsymbol{\delta}_{j}\big)g(\mathbf{L})^{\top}\boldsymbol{\delta}_{i}\boldsymbol{\delta}_{j}^{\top}\big[\mathbf{W}^{(k)}\big]^{\top}.$
@@
This completes the derivation of (A.2) and (A.3).
@@
    $\|\mathbf{X}^{(k)}\|_{F}\leq B^{k}\alpha^{k}_{\sigma}C^{k}_{g}C_{\mathbf{X}}.$ (A.5)
@@
    $\|\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\|_{F}\leq B^{K+1-k}\alpha^{K+1-k}_{\sigma}C^{K+1-k}_{g}.$ (A.6)
@@
    $\big\|\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta)\big\|_{F}\leq B^{K}\alpha_{\sigma}^{K+1}C^{K+1}_{g}C_{\mathbf{X}},$ (A.7)
@@
    Firstly, for $k=1,2,\dots,K$, since $\|\sigma(\mathbf{Z})\|_{F}\leq\alpha_{\sigma}\|\mathbf{Z}\|_{F}$ holds for any matrix $\mathbf{Z}$, we have
    $\|\mathbf{X}^{(k)}\|_{F}=\|\sigma\big(g(\mathbf{L})\mathbf{X}^{(k-1)}\mathbf{W}^{(k)}\big)\|_{F}\leq\alpha_{\sigma}\|g(\mathbf{L})\mathbf{X}^{(k-1)}\mathbf{W}^{(k)}\|_{F}.$
    Then, by applying the inequality $\|\mathbf{A}_{1}\mathbf{A}_{2}\|_{F}\leq\|\mathbf{A}_{1}\|_{2}\|\mathbf{A}_{2}\|_{F}$ twice, we obtain $\|\mathbf{X}^{(k)}\|_{F}\leq B\alpha_{\sigma}C_{g}\|\mathbf{X}^{(k-1)}\|_{F}.$ Note that $\|\mathbf{X}^{(1)}\|_{F}\leq B\alpha_{\sigma}C_{g}\|\mathbf{X}^{(0)}\|_{F}=B\alpha_{\sigma}C_{g}C_{\mathbf{X}}$, it further yields that
@@
    which completes the proof of (A.5).
@@
    To show (A.6), note that for $k=1,2,\dots,K-1$, by applying $\|\mathbf{A}_{1}\mathbf{A}_{2}\|_{F}\leq\|\mathbf{A}_{1}\|_{2}\|\mathbf{A}_{2}\|_{F}$ twice, we obtain
    $\Big\|\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\Big\|_{F}=\Big\|g(\mathbf{L})^{\top}\Big(\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k+1)}}\odot\mathbf{R}^{(k+1)}\Big)\big[\mathbf{W}^{(k+1)}\big]^{\top}\Big\|_{F}\leq\|g(\mathbf{L})\|_{2}\Big\|\Big(\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k+1)}}\odot\mathbf{R}^{(k+1)}\Big)\Big\|_{F}\|\mathbf{W}^{(k+1)}\|_{2}.$
    Since $C_{g}=\|g(\mathbf{L})\|_{2}$, $\|\mathbf{W}^{(k+1)}\|_{2}\leq B$ and the absolute value of the elements in $\mathbf{R}^{(k+1)}$ is less than $\alpha_{\sigma}$, we further have $\big\|\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\big\|_{F}\leq B\alpha_{\sigma}C_{g}\|\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k+1)}}\|_{F}$. Meanwhile, since $|\nabla\sigma\big(\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}\big)|\leq\alpha_{\sigma}$,
    $\big\|\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(K)}}\big\|_{F}=\big\|\nabla\sigma\big(\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}\big)\big[\boldsymbol{\delta}^{\top}_{\mathbf{x}}g(\mathbf{L})\big]^{\top}\mathbf{w}\big\|_{F}\leq B\alpha_{\sigma}C_{g}.$
    Therefore, for $k=1,2,\dots,K$,
@@
    This completes the proof of (A.6).
@@
    Now, let’s prove (A.7). Firstly, note that $|\nabla\sigma\big(\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}\big)|\leq\alpha_{\sigma}$, so
    $\big\|\nabla_{\mathbf{w}}f(\mathbf{x}|\theta)\big\|_{F}=\big\|\nabla\sigma\big(\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}\big)\big[g(\mathbf{L})\mathbf{X}^{(K)}\big]^{\top}\boldsymbol{\delta}_{\mathbf{x}}\big\|_{F}\leq\alpha_{\sigma}\|\mathbf{X}^{(K)}\|_{F}\|\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\|_{2}.$
    Combining (A.5) and $\|\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\|_{2}\leq C_{g}$, we have
    $\big\|\nabla_{\mathbf{w}}f(\mathbf{x}|\theta)\big\|_{F}\leq B^{K}\alpha_{\sigma}^{K+1}C^{K+1}_{g}C_{\mathbf{X}}.$
    Furthermore, for $k=1,2,\dots,K$, by applying $\|\mathbf{A}_{1}\mathbf{A}_{2}\|_{F}\leq\|\mathbf{A}_{1}\|_{2}\|\mathbf{A}_{2}\|_{F}$ twice, it yields
    $\|\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta)\|_{F}=\big\|\big[g(\mathbf{L})\mathbf{X}^{(k-1)}\big]^{\top}\Big(\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\odot\mathbf{R}^{(k)}\Big)\big\|_{F}\leq\big\|g(\mathbf{L})\big\|_{2}\big\|\mathbf{X}^{(k-1)}\big\|_{F}\big\|\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\odot\mathbf{R}^{(k)}\big\|_{F}.$
    Since the absolute value of the elements in $\mathbf{R}^{(k)}$ is less than $\alpha_{\sigma}$, we have
    $\|\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta)\|_{F}\leq\alpha_{\sigma}C_{g}\big\|\mathbf{X}^{(k-1)}\big\|_{F}\big\|\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\big\|_{F}\leq B^{K}\alpha^{K+1}_{\sigma}C^{K+1}_{g}C_{\mathbf{X}},$
    which holds by combining (A.5) and (A.6). This completes the proof of (A.7).
APPENDIX B:Proof of Lemma 2
@@
    $\|\triangle\mathbf{X}^{(k)}\|_{F}\leq B^{k-1}\alpha^{k}_{\sigma}C^{k}_{g}C_{\mathbf{X}}\Big(\sum_{j=1}^{k}\|\triangle\mathbf{W}^{(j)}\|_{2}\Big).$ (A.8)
@@
    $|f(\mathbf{x}|\theta)-f(\mathbf{x}|\theta^{\prime})|\leq B^{K}\alpha^{K+1}_{\sigma}C^{K+1}_{g}C_{\mathbf{X}}\|\triangle\theta\|_{*}.$ (A.9)
Proof: To prove (A.8), we first have that for $k=1,2,\dots,K$,
$\|\triangle\mathbf{X}^{(k)}\|_{F}=\|\mathbf{X}^{(k)}-\mathbf{X}^{(k)^{\prime}}\|_{F}=\|\sigma\big(g(\mathbf{L})\mathbf{X}^{(k-1)}\mathbf{W}^{(k)}\big)-\sigma\big(g(\mathbf{L})\mathbf{X}^{(k-1)^{\prime}}\mathbf{W}^{(k)^{\prime}}\big)\|_{F}.$
Since $\|\sigma(\mathbf{Z})\|_{F}\leq\alpha_{\sigma}\|\mathbf{Z}\|_{F}$ holds for any matrix $\mathbf{Z}$, we have
$\|\triangle\mathbf{X}^{(k)}\|_{F}\leq\alpha_{\sigma}\|g(\mathbf{L})\big(\mathbf{X}^{(k-1)}\mathbf{W}^{(k)}-\mathbf{X}^{(k-1)^{\prime}}\mathbf{W}^{(k)^{\prime}}\big)\|_{F}\leq\alpha_{\sigma}\|g(\mathbf{L})\|_{2}\cdot\|\mathbf{X}^{(k-1)}\mathbf{W}^{(k)}-\mathbf{X}^{(k-1)^{\prime}}\mathbf{W}^{(k)^{\prime}}\|_{F}.$
Note that
$\|\mathbf{X}^{(k-1)}\mathbf{W}^{(k)}-\mathbf{X}^{(k-1)^{\prime}}\mathbf{W}^{(k)^{\prime}}\|_{F}$ $\leq\|\mathbf{X}^{(k-1)}\|_{F}\|\mathbf{W}^{(k)}-\mathbf{W}^{(k)^{\prime}}\|_{2}+\|\mathbf{X}^{(k-1)}-\mathbf{X}^{(k-1)^{\prime}}\|_{F}\|\mathbf{W}^{(k)^{\prime}}\|_{2}$
$=\|\mathbf{X}^{(k-1)}\|_{F}\|\triangle\mathbf{W}^{(k)}\|_{2}+\|\triangle\mathbf{X}^{(k-1)}\|_{F}\|\mathbf{W}^{(k)^{\prime}}\|_{2}.$
Then, combining (A.5) and $\|\mathbf{W}^{(k)^{\prime}}\|_{2}\leq B$, we obtain
$\|\mathbf{X}^{(k-1)}\mathbf{W}^{(k)}-\mathbf{X}^{(k-1)^{\prime}}\mathbf{W}^{(k)^{\prime}}\|_{F}\leq B^{k-1}\alpha^{k-1}_{\sigma}C^{k-1}_{g}C_{\mathbf{X}}\|\triangle\mathbf{W}^{(k)}\|_{2}+B\|\triangle\mathbf{X}^{(k-1)}\|_{F}.$
Thus,
$\|\triangle\mathbf{X}^{(k)}\|_{F}\leq\alpha_{\sigma}\|g(\mathbf{L})\|_{2}\cdot\|\mathbf{X}^{(k-1)}\mathbf{W}^{(k)}-\mathbf{X}^{(k-1)^{\prime}}\mathbf{W}^{(k)^{\prime}}\|_{F}\leq B^{k-1}\alpha^{k}_{\sigma}C^{k}_{g}C_{\mathbf{X}}\|\triangle\mathbf{W}^{(k)}\|_{2}+B\alpha_{\sigma}C_{g}\|\triangle\mathbf{X}^{(k-1)}\|_{F}.$
Then, since $\|\triangle\mathbf{X}^{(1)}\|_{F}\leq\alpha_{\sigma}C_{g}C_{\mathbf{X}}\|\triangle\mathbf{W}^{(1)}\|_{2}$, we have